OpenStudy (anonymous):

Use factoring by difference of squares to solve the equation 4x^2 - 81 =0

OpenStudy (anonymous):

what do you square to get 4x^2? How about 81? From there, it's easy...

OpenStudy (anonymous):

4x and x is to the second power lol

OpenStudy (anonymous):

check it out... (a+b)(a-b) = a^2 + ab - ab - b^2 = a^2 - b^2 ... no middle term!!! That's why it's called a difference of squares.

OpenStudy (anonymous):

yes.. but what times itself gives you 4x^2?

OpenStudy (anonymous):

idk it's on my questions and i thnk it's 2x times 2x

OpenStudy (anonymous):

excellent!!! That's it! What about 81?

OpenStudy (anonymous):

9 ?

OpenStudy (anonymous):

OH so x = 2 or x = 9?? :D

OpenStudy (anonymous):

perfect!!! you are pretty smart! so let a = 2x and b = 9 up above...

OpenStudy (anonymous):

(2x - 9)(2x + 9)! c:

OpenStudy (anonymous):

nope, you've got to put them in the factors like I showed above. (a + b)(a - b) = (2x + 9)(2x - 9)= 0 then just set each factor = 0 and solve for x.

OpenStudy (anonymous):

you are ROCKIN!!!!

OpenStudy (anonymous):

so +2/9 and - 2/9? :D

OpenStudy (anonymous):

I mean the other way around haha, 9/2 and - 9/2? :)

OpenStudy (anonymous):

close... 2x+9 = 0 => 2x = -9 => x = -9/2 likewise for the 2x-9 factor

OpenStudy (anonymous):

dude... I mean dudette... you're so on top of this!!!

OpenStudy (anonymous):

GREAT JOB!!!

OpenStudy (anonymous):

Thank you so much!! I might need more help though with other things tho :O

OpenStudy (anonymous):

You're welcome! I'll be here and I'll come see your questions. Don't hesitate to ask!

OpenStudy (anonymous):

Thank you so much @pgpilot326

OpenStudy (anonymous):

you're welcome!